This algebra problem is about variation.
The prices of three sizes of pizza at a pizza shop are as follows.
9-inch: $8.78
12-inch: $11.78
15-inch: $14.18
You would expect that the price of a certain size of pizza would be directly proportional to its surface area. Is that the case for this pizza shop? If not, which size of pizza is the best buy?
This is really confusing. Will someone please explain and answer? Thank you in advance!

Answer by mananth(16949)

9-inch: $8.78
Area = pi*r^2
r= 4.5
Area =63.64
Price = 8.78
price / sq. inch = 8.78/63.64
=$0.138
12-inch: $11.78
r= 6.0
Area =113.14 11.78
Price = 11.78
price / sq. inch = 11.78/113.14
=0.104
15-inch: $14.18
Area = 176.79 sq. inches
Price= 14.18
price/square inch =$ 0.080
You know which is cheaper now